# A Quality-Centric Framework for Generic Deepfake Detection

### v2

@@ II-B Curriculum Learning
Figure 5: For low-quality data, we employ the FreDA module to augment their forgery quality by reducing the easy-to-recognized artifacts and enhancing their realism.
III Method
In this section, we introduce the proposed quality-centric framework. Our framework is inspired by curriculum learning that explicitly encourages the detection model to learn forgery artifacts gradually from easy to hard based on FQS. It emphasizes three key components: firstly, the evaluation of data quality; secondly, the data augmentation on low-quality data; and thirdly, the progressing use of quality-based data during training. To achieve this, we carefully design three specified modules to implement the framework: 1) FQS is proposed as a metric to assess sample hardness during model training. It consists of two main components: static scores and dynamic scores. Static score (related to data) provides initial guidance at the start of model training. The weight of the static score gradually decreases as the training progresses. Dynamic score (related to model) is continuously updated based on the model’s feedback, to better screen samples throughout the training process; 2) A FreDA module is proposed to enhance the forgery quality (“realism”) of the augmented low-quality fake samples; 3) A pacing function is presented to control the pace of presenting data from easy to hard according to FQS. Each module will be detailed subsequently.
III-A Forgery Quality Score (FQS)
@@
The hardness score plays a key role in curriculum learning as it describes the relative “hardness” of each sample in deepfake detection. In this work, we propose the FQS, which considers the dynamic model behavior and the static facial data quality.
@@
Static quality. The face recognition network learns rich facial feature representations through a large amount of external data. The features extracted through the face recognition network can capture detailed information and specific face attributes in the face image. High static quality deepfake images may have more accurate facial alignment, more realistic texture details, and more natural expression changes. Their features are more similar to the original images, causing the features extracted by the face network to be more similar to the original image features. On the contrary, low static quality deepfake images may have distortion, blur, or obvious forged details, resulting in features extracted by the face network being less similar to the original image features. For commonly used training datasets, there is a correspondence between deepfake videos and real videos. Taking advantage of such a fact, we input the fake image and its corresponding real image into a pre-trained face recognition network (Arcface), and calculate the cosine similarity of the features obtained by the model. The higher the similarity, the better the quality of the sample, and vice versa.
For an input fake image $x_{f}^{i}$, we obtain its corresponding real image $x_{r}^{i}$, input it into the face recognition network G($\cdot$), and obtain two features $g_{f}^{i}$, $g_{r}^{i}$:
$\quad\begin{cases}g_{f}^{i}=G(x_{f}^{i}),\\ g_{r}^{i}=G(x_{r}^{i}).\end{cases}$ (1)
After that, we calculate the cosine similarity between the two features as the ID similarity score $q^{i}$ of $x_{f}^{i}$:
$q^{i}=Cosine(g_{f}^{i},g_{r}^{i}).$ (2)
@@
The similarity score $q^{i}$ is used as the static quality for the forgery quality evaluation, as the more similar face-swapping pairs, the more realistic deepfake is created.
@@
Dynamic quality. Let $\mathcal{X}={(x^{i},y^{i})}_{i=1}^{N}$ be the training dataset with $N$ samples, $x^{i}$ and $y^{i}$ represent the $i$-th data and its ground-truth label, respectively. Let $f(\cdot,\theta_{t})$ be the deepfake detector model with the parameter $\theta_{t}$ at $t$-th epoch. We regard the loss (i.e., the binary cross-entropy loss denoted as $l(f(x^{i},\theta_{t-1}),y^{i})$) of $i$-th sample at $t-1$-th epoch as an indicator for the current hardness of this sample judged by the current state of the model before conducting the training step. Thus, we propose the instantaneous hardness $s_{t}^{i}$ that normalized the current loss with the learning rate $\eta_{t}$, formulated as:
$s_{t}^{i}=l(f(x^{i};\theta_{t-1}),y^{i})\cdot\eta_{max}/\eta_{t},$ (3)
where $\eta_{max}$ is the max learning rate during the training. We measure the dynamic hardness ${d}_{t}^{i}$ through a moving average of the instantaneous hardness ${s}_{t}^{i}$ over training history, defined and computed recursively as:
@@
III-B Frequency Data Augmentation (FreDA)
Samples with low (static) quality typically exhibit more pronounced distortions and artifacts compared to high-quality samples. These distortions may include severe blurring or pixelation, leading to a loss of facial detail and clarity. Additionally, low-quality samples may suffer from color inconsistencies, where certain areas of the face appear unnaturally saturated or desaturated. These inconsistencies can be attributed to improper lighting conditions or inaccuracies in the image manipulation process. Furthermore, low-quality samples often display visually evident traces of tampering, such as mismatched facial features, unnatural facial expressions, or visible blending boundaries between different facial regions, as shown in Figure 3. Overall, these factors contribute to a diminished resemblance to the original face and make low-quality samples more discernible from genuine images. According to previous work [11], we find that for fake videos of FF++, 59.44% of them are created using swapping pairs with different races or genders. Using these samples to train the model will easily make the model overfit to these obvious forgery traces, which will be harmful to the generalization of the model. However, discarding samples with low (static) quality outright would lead to a reduction in the number and diversity of training data, potentially limiting the model’s ability to generalize effectively. Therefore, it makes sense to explore an effective strategy to utilize low-quality samples for improving generalization performances.
@@ III-B Frequency Data Augmentation (FreDA)
Specifically, given a sample $x\in\mathbb{R}^{H\times W\times C}$ we perform a 2D fast Fourier transform (FFT) for each channel independently to get the corresponding frequency representation $F$ as below:
$F(u,v)=\sum_{h=0}^{H-1}\sum_{w=0}^{W-1}x(h,w)e^{-j2\pi\left(u\frac{h}{H}+v\frac{w}{W}\right)}.$ (6)
Further, we denote with $M$ a mask, whose value is zero except for the center region:
@@
where $\otimes$ denotes element-wise multiplication. Perform iFFT changes on the obtained $F_{a}$ and convert it to the RGB domain, and then obtain the FreDA image $x_{a}$. The frequency representation $F$ can be transferred to the original RGB space via an inverse FFT (iFFT), which can be formulated as:
${x}(h,w)=\frac{1}{H\cdot W}\sum_{h=0}^{H-1}\sum_{w=0}^{W-1}F(u,v)e^{j2\pi\left(u\frac{h}{H}+v\frac{w}{W}\right)}.$ (9)
Compared with $x_{f}$, $x_{a}$ has a relatively complete facial structure while retaining the traces of forgery.
@@
The proposed FreDA focuses on frequency-level augmentation and distinguishes itself from previous studies in the following ways. Compared to frequency-based detectors: FreDA aims to reduce the obvious forgery artifacts of a less realistic deepfake image and transform it into a realistic one at the frequency level, while previous frequency detectors [38, 40] mainly mine the frequency signals (e.g, high-frequency components) as the auxiliary information of the original RGB and encourage the detector to learn spatial-frequency anomalies. Compared to augmentation-based detectors: FreDA is targeted on deepfake data and designed to reduce the obvious fake artifacts of the deepfake data at the frequency level, while previous augmentation-based detectors [7, 49, 50, 10] mostly target the real data and simulate the fake blending artifacts to transform the real data into the (more realistic) pseudo-fake data.
Overall, our proposed FreDA is a novel frequency-level augmentation method that is distinct from both existing frequency detectors and augmentation-based detectors.
III-C Learning Pacing
@@
To control the learning pace of presenting data from easy to hard, we design a pacing function to determine the sample pool $\mathcal{X}^{\prime}_{t}$ of training data according to FQS. Like human education, if a teacher presents materials from easy to hard in a very short period, students will become confused and will not learn effectively. Thus, we define a pacing sequence $\mathbf{T}=[T_{0},\dots,T_{N}]$ to represent $n+1$ milestones (i.e., $n$ episodes) in total training epoch $T$. During the first $T_{0}$ epochs, we utilize all the samples in $\mathcal{X}$ for the warm-up training. After epoch $T_{0}$, we only change the size of $\mathcal{X}^{\prime}_{t}$ at every milestone $T_{n}$. Specifically, at each epoch $t$ of episode $n$, we select $k_{n}$ samples with top FQS values in $\mathcal{X}$ (i.e., hardest samples) as a hard sample pool $H_{t}$. Along with the training, we reduce the size of $H_{t}$ by $k_{n}\leftarrow\alpha_{\beta}\times k_{n-1}$ with discount factor $\alpha_{\beta}$ to make it gradually focus on harder samples. To further enlarge the diversity of the data, we also select the $E$ samples with bottom FQS values in $\mathcal{X}$ (i.e., easiest samples) as an easy sample pool $E_{t}$ and then conduct $FreDA$ on them. Finally, we get the sample pool $\mathcal{X}^{\prime}_{t}$ by mixing the $H_{t}$ and $FreDA(E_{t})$, i.e., $\mathcal{X}^{\prime}_{t}\leftarrow H_{t}\cup FreDA(E_{t})$. After that, we sample mini-batch in $\mathcal{X}^{\prime}_{t}$ and then conduct a vanilla training step to update the model parameter as $\theta_{t}$.
Our proposed strategy differs from previous CL works in two key aspects. First, our framework introduces a novel metric (FQS) to evaluate the forgery quality of deepfake samples. Additionally, we consider the model’s dynamic feedback to achieve a more comprehensive assessment of forgery quality. In contrast, previous approaches [62, 57, 59] typically only consider using a certain indicator (which may not be specifically designed for deepfake data) to measure the difficulty level of samples and ignore either the static view or the dynamic view. Second, we introduce data augmentation to improve the quality of low-quality samples (those with a lower forgery quality score) after pacing the samples. Previous approaches do not further propose improving the lower score samples but use the higher score samples only.
IV EXPERIMENTS
IV-A Experiment Settings
Datasets. In this paper, we train all models on the FaceForensics++ (FF++) [2] dataset. It contains 1,000 Pristine videos (i.e., the real sample) and 4,000 fake videos forged by four manipulation methods, i.e., Deepfakes (DF), Face2Face (F2F), FaceSwap (FS), NeuralTextures (NT). Besides, FF++ provides three quality levels in compression for these videos: raw, high-quality (HQ), and low-quality (LQ). The HQ version of FF++ is adopted by default in this paper. If any deviation from this default, it will be explicitly stated. All samples within FF++ are split into training, validation, and testing sets at the video level following the official protocol.
For cross-dataset evaluation, we utilize four recent Deepfake datasets: Celeb-DFv2 (CDF) [19] employs advanced Deepfake algorithms on celebrity videos sourced from YouTube. Deepfake Detection Challenge Preview (DFDCp) [14]: This dataset includes forgery videos featuring diverse ethnicities, ages, and genders. Deepfake Detection Challenge Public Test Set (DFDC) [20]: similar to DFDCp, this dataset contains videos with varied disturbance levels, such as compression, noise, and special effects, making it particularly challenging. WildDeepfake (Wild) [13] contains 3,805 real face sequences and 3,509 fake face sequences collected from the internet. This collection includes a variety of synthesis methods, backgrounds, and character identities. For our experiments, we utilize the standard test set, which consists of 806 sequences.
@@ IV-A Experiment Settings
Implementation details. As for pre-processing, we utilize MTCNN to detect and crop the face regions (enlarged by a factor of 1.3) from each video frame and resize them to 256 $\times$ 256. For each video, we select every ten frames to form the training dataset. We adopt Swin-Transformer-V2 Base (Swin-V2) [18] as the default backbone network, and the parameters are initialized by the weights pre-trained on the ImageNet1k. We employed the SGD optimizer with a cosine learning rate scheduler with $\eta_{max}=0.1$ (used in 5). We set 20 epoch for totally training, the pacing sequence $\mathbf{T}=[2,5,8,12,15]$ and hyper-parameters $\gamma=0.9$ (used in 4), $\alpha_{\beta}=0.9$, $E=1000$. The parameter $\alpha_{f}$ is initialized at 0.5 and undergoes a reduction by half at each milestone.
@@
Evaluation Metrics. In this work, we mainly report the area under the ROC curve (AUC) and accuracy (ACC) to compare with prior works. The video-level results are obtained by averaging predictions over each frame on an evaluated video.
IV-B Performance Comparisons
IV-B1 Intra-Dataset Comparisons
In this section, we conduct comprehensive experiments on four different face processing methods listed in the FaceForensics++ dataset. We followed the same experimental setup as previous studies [35, 43, 63], training and testing our models on LQ DF, F2F, FS, and NT respectively. Table I shows the comparison results under video-level and image-level ACC. We have selected SPSL [40], which utilizes information from the frequency domain. Additionally, we chose models that have achieved state-of-the-art (SOTA) performance in our database within the past two years for comparison, including DL [43], DIL [63], FAMM [64], DDL [35] and Yin et al. [65]. The results show that our proposed method outperforms all compared methods in all cases except for FS, which is slightly worse than DDL. DDL uses a dynamic fine-grained difference capture module and a multi-scale spatio-temporal aggregation module to capture temporal inconsistency, which may overfit to the temporal inconsistency of the known manipulation methods. Consequently, the performance of DDL on CDF and DFDC is significantly inferior to our proposed method. In addition, our method outperforms DDL by 0.6% on average, which also proves the effectiveness of our method.
TABLE I: Frame-level (top) and Vdieo-level (bottom) AUC (%) o on Faceforensics++ LQ dataset with four different manipulation methods, i.e. DeepFakes(DF), Face2Face (F2F), FaceSwap(FS), NeuralTextures(NT). The bold results are the best. * Indicates our re-implementation of the existing methods.
Methods Venue DF F2F FS NT Avg DIL [63] AAAI’22 86.78 70.85 81.59 57.64 74.22 DDL [35] TIFS’23 97.43 92.28 96.01 79.35 91.27 FAMM [64] TCSVT’23 90.00 91.00 92.75 85.50 89.81 Yin et al. [65] TMM’24 94.75 92.19 88.49 73.87 87.33 Ours - 97.84 93.58 95.56 80.32 91.83 SPSL* [40] CVPR’21 96.79 91.79 96.43 82.88 91.97 Swin-V2* [18] CVPR’22 97.50 89.64 93.57 83.93 91.16 DL [43] TIFS’22 97.25 94.46 97.13 84.63 93.37 Ours - 99.29 96.42 97.15 86.07 94.73
IV-B2 Cross-manipulation Comparisons
In real-world scenarios, predicting the methods used for face manipulation is often challenging. Therefore, it is crucial for our model to exhibit strong generalization capabilities to unseen manipulations. We conducted experiments using the FF++(HQ) dataset. We trained a model on one method from FF++ and tested it across all four methods (DF, F2F, FS, and NT). We select two state-of-the-art (SOTA) methods for comparison, including RECCE [30], CFM [47], and our baseline model Swin-V2 [18]. As shown in Table II, our proposed method can improve the average AUC by more than 10% compared with the baseline Swin-V2.
TABLE II: Image-level cross-manipulation evaluation results (AUC). Grey Background indicates intra-manipulation results. The bold results are the best.
Methods Train Venue DF F2F FS NT Avg. RECCE [30] DF CVPR’22 99.95 69.75 54.72 77.15 75.39 CFM [47] TIFS’24 99.93 77.56 54.94 75.04 76.87 Swin [18] CVPR’22 99.66 65.86 52.77 76.53 73.71 Ours - 99.41 75.32 61.57 77.81 78.53 RECCE [30] F2F CVPR’22 71.55 99.20 50.02 72.27 73.26 CFM [47] TIFS’24 81.85 99.23 60.12 70.80 78.00 Swin [18] CVPR’22 82.65 99.23 55.14 57.74 73.69 Ours - 93.94 99.13 64.86 78.43 84.09 RECCE [30] FS CVPR’22 63.05 66.21 99.72 58.07 71.76 CFM [47] TIFS’24 72.91 71.39 99.85 51.69 73.96 Swin [18] CVPR’22 54.27 65.41 99.36 58.49 69.38 Ours - 84.31 82.97 99.39 67.74 83.60 RECCE [30] NT CVPR’22 72.37 64.69 51.61 99.59 72.07 CFM [47] TIFS’24 88.31 76.78 52.56 99.24 79.22 Swin [18] CVPR’22 88.91 66.22 54.79 96.85 76.69 Ours - 94.62 83.99 56.73 97.23 83.14
In addition, our method has comprehensive improvements over the SOTA method, which shows that our method has good generalization when detecting unseen manipulation methods.
IV-B3 Cross-Datasets Comparisons
@@
The cross-datasets evaluation is still a challenging task because the unknown domain gap between the training and testing datasets can be caused by different source data, forgery methods, and/or post-processing. In this part, we evaluate generalization performance in a cross-dataset setting. Specifically, our models were trained on the FF++ and tested on CelebDF, Wild, DFDC-p, and DFDC, respectively. We compare our proposed method against twenty-three state-of-the-art DeepFake detectors in the past three years: RECCE [30], DL [43], SLADD [50], UIA-ViT [32], UCF [28], SFDG [41], NoiseDF [33], MSVT [44], FoCus [34], LSDA [9], DCL [46], CD-NeT [42], EFNB4 + SBIs [10], DDL [35], SeeABLE [48], CDIN [45], TALL [36], AUNet [51], CADDM [37], PIM [66], TDSCL [67], CFM [47] and DFFC [16]. The experimental results in terms of frame-level and video-level AUC are shown in Table III. It is evident that the proposed method lead to superior performance compared to other models in most cases, achieving the overall best results. All models are trained on the FF++ (C23) dataset for a fair comparison. We observe that our method surpasses the leading competitors in both frame-level and video-level evaluations. For instance, our approach surpasses TALL [36], which also employs Swin Transformer as its backbone network, by around 4% and 6% when testing on CDF and DFDC. The experimental results clearly highlight the significance of our method in enhancing the generalization capabilities of deepfake detectors.
@@
TABLE III: frame-level (top) and Video-level (bottom) AUC (%) of cross-datasets performances compared with SOTA methods. The results of prior methods are directly cited from the original paper and their subsequences for fair comparison. The best results are highlighted.
Methods Venue Type Test Set AUC (%) CDF Wild DFDCp DFDC RECCE [30] CVPR’22 Frame 68.71 64.31 - 69.06 DL [43] TIFS’22 Frame 70.65 - - 63.31 SLADD [50] CVPR’22 Frame 79.70 - - 77.20 UIA-ViT [32] ECCV’22 Frame 82.41 - 75.80 - UCF [28] ICCV’23 Frame 75.27 - 75.94 71.91 SFDG [41] CVPR’23 Frame 75.83 69.28 - 73.64 NoiseDF [33] AAAI’23 Frame 75.89 - - 63.89 MSVT [44] TCSVT’23 Frame 84.67 - - 73.63 FoCus [34] TIFS’24 Frame 76.13 73.31 76.62 68.42 LSDA [9] CVPR’24 Frame 83.00 - 81.50 73.60 ProgressDet [68] NeurIPS’24 Frame 84.48 - 81.16 72.40 Ours - Frame 89.18 80.74 90.94 77.65 DCL [46] AAAI22 Video 82.30 71.14 - 76.71 CD-NeT [42] ECCV’22 Video 88.50 - - 77.00 EFNB4 + SBIs [10] CVPR’22 Video 93.18 - 86.15 72.42 DDL [35] TIFS’23 Video 71.36 - - 73.08 SeeABLE [48] CVPR’23 Video 87.30 - 86.30 75.90 CDIN [45] TCSVT’23 Video 89.10 - - 78.40 TALL [36] ICCV’23 Video 90.79 - - 76.78 AUNet [51] CVPR’23 Video 92.77 - 86.16 73.82 CADDM [37] CVPR’23 Video 93.88 - - 73.85 PIM [66] TIFS’24 Video 87.34 72.28 80.69 68.04 TDSCL [67] TCSVT’24 Video 89.06 78.93 - 80.85 CFM [47] TIFS’24 Video 89.65 82.27 - 80.22 DFFC [16] ICASSP’24 Video 82.26 71.75 90.63 72.37 Ours - Video 94.75 83.77 93.41 82.04
IV-C Ablation Study

### v3

@@ 3 Method
3.1 Forgery Quality Score (FQS)
@@
The hardness score plays a key role in curriculum learning as it describes the relative “hardness" of each sample in deepfake detection. In this work, we propose the FQS, which considers the dynamic model behavior and the static facial data quality.
@@
Static quality. The face recognition network learns rich facial feature representations through a large amount of external data. The features extracted through the face recognition network can capture detailed information and specific face attributes in the face image. High-quality deepfake images may have more accurate facial alignment, more realistic texture details, and more natural expression changes. Their features are more similar to the original images, causing the features extracted by the face network to be more similar to the original image features. On the contrary, low static quality deepfake images may have distortion, blur, or obvious forged details, resulting in features extracted by the face network being less similar to the original image features. For commonly used training datasets, there is a correspondence between deepfake videos and real videos. Taking advantage of such a fact, we input the fake image and its corresponding real image into a pre-trained face recognition network (Arcface), and calculate the cosine similarity of the features obtained by the model. The higher the similarity, the better the quality of the sample, and vice versa.
@@
For an input fake image $x_{f}^{i}$, we obtain its corresponding real image $x_{r}^{i}$, input it into the face recognition network G($\cdot$), and obtain two features $g_{f}^{i}$, $g_{r}^{i}$:$g_{f}^{i}=G(x_{f}^{i})$, $g_{r}^{i}=G(x_{r}^{i})$. After that, we calculate the cosine similarity between the two features as the ID similarity score $q^{i}$ of $x_{f}^{i}$: $q^{i}=Cosine(g_{f}^{i},g_{r}^{i})$. The similarity score $q^{i}$ is used as the static quality for the forgery quality evaluation, as the more similar face-swapping pairs, the more realistic the deepfake is created.
@@
Dynamic quality. Let $\mathcal{X}={(x^{i},y^{i})}_{i=1}^{N}$ be the training dataset with $N$ samples, $x^{i}$ and $y^{i}$ represent the $i$-th data and its ground-truth label, respectively. Let $f(\cdot,\theta_{t})$ be the deepfake detector model with the parameter $\theta_{t}$ at $t$-th epoch. We regard the loss (i.e., the binary cross-entropy loss denoted as $l(f(x^{i},\theta_{t-1}),y^{i})$) of $i$-th sample at $t-1$-th epoch as an indicator for the current hardness of this sample judged by the current state of the model before conducting the training step. Thus, we propose the instantaneous hardness $s_{t}^{i}$ that normalized the current loss with the learning rate $\eta_{t}$, formulated as: $s_{t}^{i}=l(f(x^{i};\theta_{t-1}),y^{i})\cdot\eta_{max}/\eta_{t},$ where $\eta_{max}$ is the max learning rate during the training. We measure the dynamic hardness ${d}_{t}^{i}$ through a moving average of the instantaneous hardness ${s}_{t}^{i}$ over training history, defined and computed recursively as:
@@
3.2 Frequency Data Augmentation (FreDA)
Figure 3: For low-quality data, we employ the FreDA module to augment their forgery quality by reducing the easily recognizable artifacts and enhancing their realism.
@@ 3.2 Frequency Data Augmentation (FreDA)
Specifically, given a sample $x\in\mathbb{R}^{H\times W\times C}$ we perform a 2D fast Fourier transform (FFT) for each channel independently to get the corresponding frequency representation $F$. Further, we denote with $M$ a mask, whose value is zero except for the center region:
@@
where $\otimes$ denotes element-wise multiplication. Perform iFFT changes on the obtained $F_{a}$ and convert it to the RGB domain, and then obtain the FreDA image $x_{a}$. The frequency representation $F$ can be transferred to the original RGB space via an inverse FFT (iFFT). Compared with $x_{f}$, $x_{a}$ has a relatively complete facial structure while retaining the traces of forgery.
@@
The proposed FreDA focuses on frequency-level augmentation and distinguishes itself from previous studies in the following ways. Compared to frequency-based detectors: FreDA aims to reduce the obvious forgery artifacts of a less realistic deepfake image and transform it into a realistic one at the frequency level, while previous frequency detectors [38, 33] mainly mine the frequency signals (e.g, high-frequency components) as the auxiliary information of the original RGB and encourage the detector to learn spatial-frequency anomalies. Compared to augmentation-based detectors: FreDA is targeted on deepfake data and designed to reduce the obvious fake artifacts of the deepfake data at the frequency level, while previous augmentation-based detectors [28, 63, 6, 40] mostly target the real data and simulate the fake blending artifacts to transform the real data into the (more realistic) pseudo-fake data. Overall, our proposed FreDA is a novel frequency-level augmentation method that is distinct from both existing frequency detectors and augmentation-based detectors.
3.3 Learning Pacing
@@
To control the learning pace of presenting data from easy to hard, we design a pacing function to determine the sample pool $\mathcal{X}^{\prime}_{t}$ of training data according to FQS. Like human education, if a teacher presents materials from easy to hard in a very short period, students will become confused and will not learn effectively. Thus, we define a pacing sequence $\mathbf{T}=[T_{0},\dots,T_{N}]$ to represent $n+1$ milestones (i.e., $n$ episodes) in total training epoch $T$. During the first $T_{0}$ epochs, we utilize all the samples in $\mathcal{X}$ for the warm-up training. After epoch $T_{0}$, we only change the size of $\mathcal{X}^{\prime}_{t}$ at every milestone $T_{n}$. Specifically, at each epoch $t$ of episode $n$, we select $k_{n}$ samples with top FQS values in $\mathcal{X}$ (i.e., hardest samples) as a hard sample pool $H_{t}$. Along with the training, we reduce the size of $H_{t}$ by $k_{n}\leftarrow\alpha_{\beta}\times k_{n-1}$ with discount factor $\alpha_{\beta}$ to make it gradually focus on harder samples. To further enlarge the diversity of the data, we also select the $E$ samples with bottom FQS values in $\mathcal{X}$ (i.e., easiest samples) as an easy sample pool $E_{t}$ and then conduct $FreDA$ on them. Then, we get the sample pool $\mathcal{X}^{\prime}_{t}$ by mixing the $H_{t}$ and $FreDA(E_{t})$, i.e., $\mathcal{X}^{\prime}_{t}\leftarrow H_{t}\cup FreDA(E_{t})$. Finally, we sample a mini-batch in $\mathcal{X}^{\prime}_{t}$ and then conduct a vanilla training step to update the model parameter as $\theta_{t}$.
4 EXPERIMENTS
4.1 Experiment Settings
Datasets. In this paper, we train all models on the FaceForensics++ (FF++) [39] dataset. To demonstrate the performance of our proposed method in cross-dataset settings, four additional datasets are adopted, i.e., Celeb-DFv2 (CDF) [31], Deepfake Detection Challenge Preview (DFDC-p) [12], Deepfake Detection Challenge Public Test Set (DFDC) [11], and WildDeepfake (Wild) [66]. For cross-manipulation evaluation, we utilize two Deepfake datasets: FF++ (train a model on one method from FF++ and test it across all four methods (DF, F2F, FS, and NT)) and DF40 (train a model on FF++ and test it across the methods in DF40) .
@@ 4.1 Experiment Settings
Implementation details. As for pre-processing, we utilize MTCNN to detect and crop the face regions (enlarged by a factor of 1.3) from each video frame and resize them to 256 $\times$ 256. For each video, we select every 10 frames to form the training dataset. We adopt Swin-Transformer-V2 Base (Swin-V2) [34] as the default backbone network, and the parameters are initialized by the weights pre-trained on the ImageNet1k. We employed the SGD optimizer with a cosine learning rate scheduler with $\eta_{max}=0.1$ (used in 2). We set 20 epochs for totally training, the pacing sequence $\mathbf{T}=[2,5,8,12,15],$ and hyper-parameters $\gamma=0.9$ (used in 1), $\alpha_{\beta}=0.9$, $E=1000$. The parameter $\alpha_{f}$ is initialized at 0.5 and undergoes a reduction by half at each milestone.
@@
Evaluation Metrics. In this work, we mainly report the area under the ROC curve (AUC) to compare with prior works. The video-level results are obtained by averaging predictions over each frame on an evaluated video.
4.2 Performance Comparisons
4.2.1 Cross-Datasets Comparisons
@@
The cross-dataset evaluation is still a challenging task because the unknown domain gap between the training and testing datasets can be caused by different source data, forgery methods, and/or post-processing. In this part, we evaluate generalization performance in a cross-dataset setting. Specifically, our models were trained on the FF++ and tested on CelebDF, Wild, DFDC-p, and DFDC, respectively. We compare our proposed method against twenty-three state-of-the-art DeepFake detectors in the past three years: RECCE [4], SLADD [6], UIA-ViT [65], PEL [17], UCF [57], SFDG [50], NoiseDF [49], FoCus [45], LSDA [53], ProgressDet [7], DCL [43], CD-NeT [41], EFNB4 + SBIs [40], CFM [35], SeeABLE [25], TALL [52], AUNet [2], PIM [19], DFFC [42] and FreqBlender [64]. The experimental results in terms of frame-level and video-level AUC are shown in Table 1.
(a) The performance of cross-manipulation methods on the FF++.
(b) The performance of cross-manipulation methods on the DF40.
Figure 4: Performance comparison of cross-manipulation with other SOTA detectors under the latest DF40 dataset.
It is evident that the proposed method lead to superior performance compared to other models in most cases, achieving the overall best results. All models are trained on the FF++ (C23) dataset for a fair comparison. We observe that our method surpasses the leading competitors in both frame-level and video-level evaluations. For instance, our approach surpasses TALL [52], which also employs Swin Transformer as its backbone network, by around 4% and 6% when testing on CDF and DFDC. The experimental results clearly highlight the significance of our method in enhancing the generalization capabilities of deepfake detectors.
@@
Table 1: Frame-level (left) and Video-level (right) AUC (%) of cross-datasets performances compared with SOTA methods. The top two results are highlighted, with the best in bold and the second-best underlined. ‘*’ indicates our reproductions.
Frame-Level AUC Video-Level AUC
Method Venues CDF Wild DFDC-p DFDC Avg Method Venues CDF Wild DFDC-p DFDC Avg
RECCE CVPR 2022 82.30 75.63 74.91 69.06 75.48 DCL AAAI 2022 82.30 71.14 - 76.71 76.72
SLADD CVPR 2022 83.70 69.02 75.60 77.20 76.39 CD-Net ECCV 2022 88.50 77.00
UIA-ViT ECCV 2022 82.41 75.80 EFNB4 + SBIs CVPR 2022 93.18 86.15 72.42 83.92
PEL AAAI 2022 69.18 67.39 63.31 66.63 CFM TIFS 2023 89.62 82.27 80.22 84.04
UCF ICCV 2023 75.27 77.40 75.94 71.91 75.13 SeeABLE CVPR 2023 87.30 86.30 75.90 83.17
SFDG CVPR 2023 75.83 69.28 73.64 72.92 TALL ICCV 2023 90.79 76.78
NoiseDF AAAI 2023 75.89 63.89 AUNet CVPR 2023 92.77 86.16 73.82 84.25
FoCus TIFS 2024 76.13 73.31 76.62 68.42 73.62 PIM TIFS 2024 87.34 72.28 80.69 68.04 77.09
LSDA CVPR 2024 83.00 81.50 73.60 79.37 DFFC ICASSP 2024 82.26 71.75 90.63 72.37 79.25
ProgressDet NeurIPS 2024 84.48 81.16 72.40 79.35 Freq NeurIPS 2024 94.59 87.56 74.59 85.58
Ours 89.18 80.74 90.94 77.65 84.63 Ours 94.75 83.77 93.41 82.04 88.49
(+4.70%) (+7.43%) (+9.44%) (+0.45%) (+5.26%) (+0.16%) (+1.50%) (+2.78%) (+1.82%) (+2.91%)
4.2.2 Cross-manipulation Comparisons
In real-world scenarios, predicting the methods used for face manipulation is often challenging. Therefore, it is crucial for our model to exhibit strong generalization capabilities to unseen manipulations. We conducted experiments using the FF++(HQ) dataset. We trained a model on one method from FF++ and tested it across all four methods (DF, F2F, FS, and NT). We select three state-of-the-art (SOTA) methods for comparison, including RECCE [4], DCL [43], CFM [35], and our baseline model Swin-V2 [34]. As shown in Fig 4(a), our proposed method can improve the average AUC by more than 10% compared with the baseline Swin-V2.
In addition, we also trained on FF++, and tested on six datasets in DF40, namely Inswap, Uniface, Blendface, Smiswap, FSGAN, and Mobileswap, the results are shown in Fig 4(b). The comparative methods include advanced ones such as SLADD [6], RECCE [4], UCF [57], LSDA [53] and ProgressDet [7]. Our method has improvements over the SOTA method, which shows that our method has good generalization when detecting unseen manipulation methods.
4.3 Ablation Study